OpenStudy (anonymous):

x + 3y = 5 -x + 6y = 4

OpenStudy (anonymous):

I x+ 3y = 5 II -x +6y = 4 if we multply the I with (-2) we will get -2x -6y =-10; we sum the two equations ; -2x-6y = -10; -x +6y = 4 -3x = -6 x = 2 and y = 1;

OpenStudy (anonymous):

wow ur smart!

OpenStudy (zzr0ck3r):

you could have added them as they are then 9y=9

OpenStudy (anonymous):

yes zzr0ck3r you have easy way ; to solve